OpenStudy (anonymous):

what do i do to solve this and give me the answer please:] 11m+13=m+23

OpenStudy (anonymous):

m=1

OpenStudy (aroub):

11m-m=-13+23 10m=10 m=1

OpenStudy (anonymous):

\[The\ solution\ is\ m = 1\]

OpenStudy (moongazer):

aroub solved it 11m-m=23-13 10m=10 (divide both sides by 10 to cancel out the 10 in 10m so it will become m=?) m=1 this is what you call transposing

OpenStudy (moongazer):

just transfer the numbers that have common variables and when you move to another side you will change the sign

OpenStudy (moongazer):

do you know now how to do it?